Survey of Study Habits and Attitudes (SSHA) is a psychological test that measures the motivational attitude toward school, and study habits of students. Scores range from 0 to 200. The mean for Canadian college students is about 115, and the standard deviation is about 30. A teacher who suspects this population average to be incorrect gives the SSHA to 25 students who are at least 30 years of age. The sample average score is 127.8. Calculate the p-value.

Select one:

a. 0.03

b. 0.3

c. 0.003

d. 0.5
Answer #1

given data l=115 7 = 30 m=25 =127.8 claim: 4 & 115 Ho: M = 115 Mail & 115 3-4 Z = Test Stretistics, olun = 127.8-115 30 125 =